# Birch SGD: A Tree Graph Framework for Local and Asynchronous SGD Methods

### v1

@@ 2.1 Main theoretical result on convergence rates
For all $y\in V,$ the representation $\textnormal{repr}(y)$ is the multiset of stochastic gradients applied to $w^{0}$ to get $y.$ In other words, there exist $\{(w_{1},\xi_{1}),\dots,(w_{p},\xi_{p})\}=:\textnormal{repr}(y)$ for some $p\geq 0$ such that $y=w^{0}-\gamma\sum_{i=1}^{p}\nabla f(w_{i},\xi_{i}).$
@@
Condition 1: For all $k\geq 0,$ $\xi^{k}$ is statistically independent of $\{(x^{i},z^{i},\xi^{i})\}_{i=0}^{k-1}.$
@@
Condition 1. The first condition condition requires that $\xi^{k}$ is independent of $\{(x^{i},z^{i},\xi^{i})\}_{i=0}^{k-1},$ which is a weak assumption. In Vanilla SGD, where $x^{k+1}=x^{k}-\gamma\nabla f(x^{k};\xi^{k})$, it is standard to assume that each $\xi^{k}$ is an independent sample. Our condition generalizes this to other SGD variants. It guarantees that the stochastic gradient $\nabla f(\cdot;\xi^{k})$ is not used in computing $x^{k}$ or $z^{k}$. Notably, this remains true even in methods like Local SGD, where gradients may be reused.
@@
Vanilla SGD (Section E.1). For instance, consider the simplest method, the classical stochastic gradient descent (Vanilla SGD) method: $w^{k+1}=w^{k}-\gamma\nabla f(w^{k};\eta^{k}),$ where $w^{0}$ is a starting point and are $\{\eta^{k}\}$ are i.i.d. random variables. Taking $x^{k}=z^{k}=w^{k}$ and $\xi^{k}=\eta^{k}$ for all $k\geq 0.$ Clearly, all conditions of Theorem 2.4 are satisfied: $\xi^{k}$ is independent of $\{(x^{i},z^{i},\xi^{i})\}_{i=0}^{k-1},$ $\textnormal{repr}(x^{k})=\textnormal{repr}(z^{k})$ for all $k\geq 0,$ and $R=0.$ We get the iteration rate $\mathcal{O}\left(\nicefrac{{L\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{\varepsilon^{2}}}\right).$ The corresponding tree is in Figure 13.
@@ C.5 Parameters of the experiments
Next, using the variance decomposition equality, we get
${\mathbb{E}}\left[\left\|\nabla f(x^{k})-\nabla f(z^{k})\right\|^{2}\right]\leq L^{2}\gamma^{2}{\mathbb{E}}\left[\left\|\sum_{j\in\bar{S}^{k}}\nabla f(z^{j})\right\|^{2}\right]+L^{2}\gamma^{2}{\mathbb{E}}\left[\left\|\sum_{j\in\bar{S}^{k}}(\nabla f(z^{j};\xi^{j})-\nabla f(z^{j}))\right\|^{2}\right].$
Since $\xi^{j}$ is statistically independent of $\{(x^{i},z^{i},\xi^{i})\}_{i=0}^{j-1}$ for all $j\in\bar{S}^{k}$ (Condition 1) and using Assumption 1.3,
${\mathbb{E}}\left[\left\|\nabla f(x^{k})-\nabla f(z^{k})\right\|^{2}\right]$ $\leq L^{2}\gamma^{2}{\mathbb{E}}\left[\left\|\sum_{j\in\bar{S}^{k}}\nabla f(z^{j})\right\|^{2}\right]+L^{2}\gamma^{2}\left|\bar{S}^{k}\right|\sigma^{2}$
@@
${\mathbb{E}}\left[\left\|\nabla f(x^{k})-\nabla f(z^{k})\right\|^{2}\right]$ $\leq L^{2}\gamma^{2}R\sum_{j=k-R}^{k-1}{\mathbb{E}}\left[\left\|\nabla f(z^{j})\right\|^{2}\right]+L^{2}\gamma^{2}R\sigma^{2}.$
@@
$\quad+\frac{\gamma}{2}\left(L^{2}\gamma^{2}R\sum_{j=k-R}^{k-1}{\mathbb{E}}\left[\left\|\nabla f(z^{j})\right\|^{2}\right]+L^{2}\gamma^{2}R\sigma^{2}\right)$
@@
$\quad+\frac{L^{2}\gamma^{3}R}{2}\sum_{j=k-R}^{k-1}{\mathbb{E}}\left[\left\|\nabla f(z^{j})\right\|^{2}\right]$ (10)
because $\gamma\leq\frac{1}{LR}.$ Note that $\sum_{k=0}^{K-1}\sum_{j=k-R}^{k-1}{\mathbb{E}}\left[\left\|\nabla f(z^{j})\right\|^{2}\right]\leq R\sum_{k=0}^{K-1}{\mathbb{E}}\left[\left\|\nabla f(z^{j})\right\|^{2}\right].$ Thus, summing (10) for $k=0,\dots,K-1$ and substituting $f^{*},$
@@
$\quad+\frac{L^{2}\gamma^{3}R^{2}}{2}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\left\|\nabla f(z^{j})\right\|^{2}\right]$
@@ E.2 Rennala SGD
Clearly, $\{x^{k}\}_{k\geq 0}$ is a main branch and $\{(z^{k},\xi^{k})\}_{k\geq 0}$ is the corresponding sequence by the construction in (12). Moreover, $\xi^{k}$ is independent of $\{(x^{i},z^{i},\xi^{i})\}_{i=0}^{k-1}$ in (12) because $\{\xi^{i}\}$ are i.i.d. (Condition 1 is satisfied). Next, notice that
@@ Appendix G Total Time Complexities of Algorithms under $(h,\tau)$-Fixed Computation Model
Note that it is possible to assume that each worker has its own communication time bound $\tau_{i}$ and computation time bound $h_{i}$ and consider $(h_{i},\tau_{i})$-fixed computation model (Tyurin et al.,, 2024). However, for simplicity, we assume $\tau_{i}=\tau$ and $h_{i}=h$ for all $i\in[n].$

### v2

@@ 2.1 Main theoretical result on convergence rates
For all $y\in V,$ the representation $\textnormal{repr}(y)$ is the multiset of stochastic gradients applied to $w^{0}$ to get $y.$ In other words, there exist $\{(m^{1},\kappa^{1}),\dots,(m^{p},\kappa^{p})\}=:\textnormal{repr}(y)$ for some $p\geq 0$ such that $y=w^{0}-\gamma\sum_{j=1}^{p}\nabla f(m^{j},\kappa^{j}).$
@@
Condition 1: For all $k\geq 0,$ $\xi^{k}$ is statistically independent of $\{(x^{i+1},z^{i+1},\xi^{i})\}_{i=0}^{k-1}.$
@@
Condition 1. The first condition condition requires that $\xi^{k}$ is independent of $\{(x^{i+1},z^{i+1},\xi^{i})\}_{i=0}^{k-1},$ which is a weak assumption. In Vanilla SGD, where $x^{k+1}=x^{k}-\gamma\nabla f(x^{k};\xi^{k})$, it is standard to assume that each $\xi^{k}$ is an independent sample. Our condition generalizes this to other SGD variants. It guarantees that the stochastic gradient $\nabla f(\cdot;\xi^{k})$ is not used in computing $x^{k}$ or $z^{k}$. Notably, this remains true even in methods like Local SGD, where gradients may be reused.
@@
Vanilla SGD (Section E.1). For instance, consider the simplest method, the classical stochastic gradient descent (Vanilla SGD) method: $w^{k+1}=w^{k}-\gamma\nabla f(w^{k};\eta^{k}),$ where $w^{0}$ is a starting point and are $\{\eta^{k}\}$ are i.i.d. random variables. Taking $x^{k}=z^{k}=w^{k}$ and $\xi^{k}=\eta^{k}$ for all $k\geq 0.$ Clearly, all conditions of Theorem 2.4 are satisfied: $\xi^{k}$ is independent of $\{(x^{i+1},z^{i+1},\xi^{i})\}_{i=0}^{k-1},$ $\textnormal{repr}(x^{k})=\textnormal{repr}(z^{k})$ for all $k\geq 0,$ and $R=0.$ We get the iteration rate $\mathcal{O}\left(\nicefrac{{L\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{\varepsilon^{2}}}\right).$ The corresponding tree is in Figure 13.
@@ C.5 Parameters of the experiments
Next, using Young’s inequality $\left\|x+y\right\|^{2}\leq 2\left\|x\right\|^{2}+2\left\|y\right\|^{2}$ for all $x,y\in\mathbb{R}^{d}$, we get
${\mathbb{E}}\left[\left\|\nabla f(x^{k})-\nabla f(z^{k})\right\|^{2}\right]\leq 2L^{2}\gamma^{2}{\mathbb{E}}\left[\left\|\sum_{j\in\bar{S}^{k}}\nabla f(z^{j})\right\|^{2}\right]+2L^{2}\gamma^{2}{\mathbb{E}}\left[\left\|\sum_{j\in\bar{S}^{k}}(\nabla f(z^{j};\xi^{j})-\nabla f(z^{j}))\right\|^{2}\right].$
Since $\xi^{j}$ is statistically independent of $\{(x^{i+1},z^{i+1},\xi^{i})\}_{i=0}^{j-1}$ for all $j\in\bar{S}^{k}$ (Condition 1) and using Assumption 1.3,
${\mathbb{E}}\left[\left\|\nabla f(x^{k})-\nabla f(z^{k})\right\|^{2}\right]$ $\leq 2L^{2}\gamma^{2}{\mathbb{E}}\left[\left\|\sum_{j\in\bar{S}^{k}}\nabla f(z^{j})\right\|^{2}\right]+2L^{2}\gamma^{2}\left|\bar{S}^{k}\right|\sigma^{2}$
@@
${\mathbb{E}}\left[\left\|\nabla f(x^{k})-\nabla f(z^{k})\right\|^{2}\right]$ $\leq 2L^{2}\gamma^{2}R\sum_{j=k-R}^{k-1}{\mathbb{E}}\left[\left\|\nabla f(z^{j})\right\|^{2}\right]+2L^{2}\gamma^{2}R\sigma^{2}.$
@@
$\quad+\frac{\gamma}{2}\left(2L^{2}\gamma^{2}R\sum_{j=k-R}^{k-1}{\mathbb{E}}\left[\left\|\nabla f(z^{j})\right\|^{2}\right]+2L^{2}\gamma^{2}R\sigma^{2}\right)$
@@
$\quad+L^{2}\gamma^{3}R\sum_{j=k-R}^{k-1}{\mathbb{E}}\left[\left\|\nabla f(z^{j})\right\|^{2}\right]$ (10)
because $\gamma\leq\frac{1}{2RL}.$ Note that $\sum_{k=0}^{K-1}\sum_{j=k-R}^{k-1}{\mathbb{E}}\left[\left\|\nabla f(z^{j})\right\|^{2}\right]\leq R\sum_{k=0}^{K-1}{\mathbb{E}}\left[\left\|\nabla f(z^{k})\right\|^{2}\right].$ Thus, summing (10) for $k=0,\dots,K-1$ and substituting $f^{*},$
@@
$\quad+L^{2}\gamma^{3}R^{2}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\left\|\nabla f(z^{k})\right\|^{2}\right]$
@@ E.2 Rennala SGD
Clearly, $\{x^{k}\}_{k\geq 0}$ is a main branch and $\{(z^{k},\xi^{k})\}_{k\geq 0}$ is the corresponding sequence by the construction in (12). Moreover, $\xi^{k}$ is independent of $\{(x^{i+1},z^{i+1},\xi^{i})\}_{i=0}^{k-1}$ in (12) because $\{\xi^{i}\}$ are i.i.d. (Condition 1 is satisfied). Next, notice that
@@ Appendix G Total Time Complexities of Algorithms under $(h,\tau)$-Fixed Computation Model
Note that it is possible to assume that each worker has its own communication time bound $\tau_{i}$ and computation time bound $h_{i}$ and consider $(h_{i},\tau_{i})$-fixed computation model (Tyurin et al.,, 2024). However, for simplicity, we assume $\tau_{i}=\tau$ and $h_{i}=h$ for all $i\in[n].$ See Section H for a more general case $(h_{i},\tau_{i})$-fixed computation model.